OpenStudy (anonymous):

write the linear equation y+3=(2/5) (x+9)

OpenStudy (anonymous):

y+3-3=(2/5) (x+9)-3 y=2x/5 +2*9/5-3/1*5/5= 2x/5 +18/5-15/5= 2x/5 +3/5 or 1/5(2x+3)

OpenStudy (anonymous):

y=2x/5 +3/5 or y=1/5(2x+3)